# UniHDA: Towards Universal Hybrid Domain Adaptation of Image Generators

### 23 Jan 2024 v1

@@ 4.4 Image-text Hybrid Domain Adaptation
Fig. 6 shows the results of image-text adaptation. As depicted in Sec. 4.2, NADA is susceptible to overfitting, which retains poor cross-domain consistency. In contrast, UniHDA (Ours) achieves robust cross-domain consistency and generates diverse images in all scenarios. Furthermore, we showcase the results of image-text hybrid domain adaptation on dogs [3], and churches [37] in Fig. 7. We can clearly observe that it achieve desirable results across various target domains. More results are included in Appendix.
@@ 4.5 Comparison with Existing Methods
In  Sec. 4.2 and  Sec. 4.3 , we find that while DE [18] and NADA [5] compromise cross-domain consistency, UniHDA (Ours) retains strong consistency and successfully captures the attributes of target domains. Besides, UniHDA also surpasses them in terms of efficiency, e.g., model size and training time as shown in Tab. 3. NADA trains a separate generative model per domain and interpolates their parameters in test-time, which necessitates multiple times the model size and training time. Although DE avoids cross-model interpolation, it heavily relies on the large source dataset (e.g., FFHQ [8]) for regularization during training process, resulting in a significant increase in training time. In contrast, UniHDA circumvents these issues, which enables the completion of the adaptation within single generator and only two minutes.
@@
Additionally, DE relies on the semantic latent space of the generator (e.g., StyleGAN [8] and DiffAE [23]) for hybrid domain adaptation, limiting its applicability to a broader range of generators. Conversely, UniHDA is not constrained by the type of generators, allowing for its broader application across various generators.
4.6 Generalization on Diffusion Models
@@
In this section, we demonstrate that UniHDA is agnostic to the type of generative models and can easily generalize to diffusion models. Specifically, we apply UniHDA on DiffusionCLIP [11] and replace the training objective of DiffusionCLIP with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$. As shown in Fig. 8, the results work as well as those on StyleGAN2, which integrate the characteristics from multiple target domains and maintain robust consistency with the source domain. This verifies desirable generalization ability and versatility of UniHDA.
4.7 Effective of CSS Loss
@@
We conduct the ablation study to evaluate the effects of our proposed Cross-domain Spatial Structure (CSS) loss on single domains. As shown in Fig. 9, the results without $\mathcal{L}_{\text{CSS}}$ suffer from overfitting and have very limited cross-domain consistency, e.g., distorted backgrounds in Row 1 and 3. Benefited from CSS, the generated images maintain consistency with the source images in terms of spatial structure, thereby inheriting the diversity from the source domain. More results are included in Appendix.
@@ 5 Conclusion & Limitation
In this paper, we propose UniHDA, a unified and versatile framework for multi-modal hybrid domain adaptation. To enable multiple modalities, we leverage CLIP encoder to project multi-modal references into a unified embedding space. For hybrid domain, we demonstrate the compositional capabilities of direction vectors in CLIP’s embedding space and linearly interpolate direction vectors of multiple target domains. In addition, we propose a new cross-domain spatial structure loss to improve cross-domain consistency by preserving intricate spatial information between the source and target generators. We believe that our work is an important step towards generative domain adaptation since we have demonstrated that the source generator can be effectively adapted to a hybrid domain with multi-modal references.
@@
While UniHDA effectively realizes multi-modal hybrid domain adaptation, it also has the limitation. To encode both image and text into a shared embedding space, we utilize pre-trained CLIP during training time, which might bring potential bias for some domains. Nevertheless, we believe that the exploration of the novel task are significant for future work and solutions could be integrated into UniHDA to eliminate the bias.
Supplementary Material
Figure 10: More qualitative results to verify the effectiveness of our proposed $\mathcal{L}_{\text{CSS}}$.
Figure 12: Hybrid domain adaptation from AFHQ cat to incompatible domains, i.e. lion, rabbit, and tiger.
Figure 13: Hybrid domain adaptation from AFHQ dog to incompatible domains, i.e. lion, rabbit, and tiger.
Figure 14: Effect of different pre-trained image encoders for CSS on image-image hybrid domain adaptation.
Figure 15: Effect of different pre-trained image encoders for CSS on text-text hybrid domain adaptation.
Figure 16: Effect of different pre-trained image encoders for CSS on image-text hybrid domain adaptation.
Figure 17: More results of UniHDA with DiffusionCLIP.
Figure 18: More results of image-image hybrid domain adaptation. The source image is in the top-left corner, and the first row and column consist of training images.
Figure 19: More results of text-text hybrid domain adaptation. The source image is in the top-left corner, and the first row and column consist of text prompts.
Figure 20: More results of image-text hybrid domain adaptation. The source image is in the top-left corner. The first row and column consist of training images and text prompts respectively.
Figure 21: More results of image-text hybrid domain adaptation on AFHQ dog.
6 Appendix
$\mathcal{L}_{\text{CSS}}$ Elena Dog Church
0.569 0.429 0.617
0.748 0.671 0.752
Table 4: Quantitative study of our $\mathcal{L}_{\text{CSS}}$, which significantly improves cross-domain consistency and achieves better SCS score.
6.1 More Ablation of CSS Loss
As depicted in Sec. 4.7, our proposed $\mathcal{L}_{\text{CSS}}$ significantly alleviates overfitting and improves cross-domain consistency. We conduct the quantitative study in Tab. 4, which aligns with the earlier qualitative results in Fig. 9. Results with $\mathcal{L}_{\text{CSS}}$ achieve better SCS score, indicating that they maintain stronger consistency with the source domain. Additionally, we show more qualitative results in Fig. 10 to verify the effectiveness of UniHDA.
6.2 Hybrid Domain Adaptation of 3D Generator
To verify the versatility of UniHDA, We apply it on the popular 3D-aware image generation method, EG3D [2]. Specifically, we replace the discrimination loss with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$ for hybrid domain adaptation. As shown in Fig. 11, the results effectively integrate the attributes and preserve the characters and poses of source domain.
6.3 Results of Incompatible Domain Adaptation
To verify the generalization ability of UniHDA, we additionally conduct experiments for hybrid domain adaptation on incompatible domains, i.e. from cat to rabbit. As shown in Fig. 12 and Fig. 13, we start from AFHQ cat and dog to incompatible domains respectively. We can observe that the results effectively integrate the attributes of the corresponding domain and maintain robust consistency with source domain.
6.4 Effect of Encoder for CSS
We conduct experiments on pre-trained ViT, MViTv2, and Dinov2 to explore the impact of different image encoders for CSS. As shown in Fig. 14, Fig. 15, and Fig. 16, we showcase the results of image-image, text-text, and image-text respectively. We can observe that all of them improve the consistency with source domain compared with the baseline approach. Furthermore, they exhibit a similar qualitative style, which demonstrates that our CSS is agnostic to different pre-trained image encoders.
6.5 More Results for DiffusionCLIP
To demonstrate the versatility of UniHDA, we apply it on DiffusionCLIP in Sec. 4.6. As shown in Fig. 17, we showcase more results for DiffusionCLIP including image-image, text-text, and image-text. All results achieve hybrid domain adaptation and preserve strong corss-domain consistency.
6.6 More Qualitative Results
We apply UniHDA to adapt the generator on FFHQ to more hybrid domains, i.e., text-text, image-image, and image-text. As shown in Fig. 18, Fig. 19, and Fig. 20, UniHDA successfully generates images with integrated characteristics from multiple target domains and maintains robust consistency with the source domain. Besides, we showcase more results of hybrid domain adaptation from AFHQ dog in Fig. 21.
6.7 Implement Details
Following the setting of previous generative domain adaptation methods [5, 18], we utilize the batch size of 4 and ADAM Optimizer with a learning rate of 0.002 for all experiments during training. A training session typically requires 300 iterations in 2 minutes, which significantly reduces training time compared with adversarial methods for generative domain adaptation. Note that we conduct all experiments on a single NVIDIA RTX 4090 GPU. The code will be open source.
For experiments on FFHQ, we generate images with 1024 × 1024 resolution. As for AFHQ-Dog and LSUN-Church, we operate on 512 × 512 and 256 × 256 resolution images respectively.

### 15 Mar 2024 v2

@@ 4.4 Image-text Hybrid Domain Adaptation
Fig. 7 shows the results of image-text adaptation, including FFHQ [12], AFHQ-Dog [5], and LSUN -Church [50]. As depicted in Sec. 4.2, NADA is susceptible to overfitting, which retains poor cross-domain consistency. Besides, we interpolate NADA’s parameters with MTG[55] and DiFa [51], which alleviates overfitting to some extent. However, they can’t accurately capture the attributes of hybrid target domain and still fail to maintain good consistency. In contrast, UniHDA well captures the attributes and achieves robust consistency in all scenarios.
As shown in Tab. 4, we also compare UniHDA with the baselines quantitatively. Consistent with qualitative results in Fig. 7, ours clearly outperforms the baselines. Besides, we conduct user study in Tab. 4. For each case, we generate 1000 samples and randomly assign 200 samples to 30 users. The results indicate that UniHDA surpasses NADA in terms of fidelity, diversity and reference correspondence.
@@ 4.5 Comparison with Existing Methods
In addition to surpassing baselines for generation quality, UniHDA also surpasses them in terms of efficiency, e.g., model size and training time as shown in Tab. 5. NADA, MTG, and DiFa trains a separate generative model per domain and interpolates their parameters in test-time, which necessitates multiple times the model size and training time. Although DE avoids cross-model interpolation, it heavily relies on the large source dataset (e.g., FFHQ [12]) for regularization during training process, resulting in a significant increase in training time. In contrast, UniHDA circumvents these issues, which enables the completion of the adaptation within single generator in only two minutes.
@@
Furthermore, DE relies on the semantic latent space of the generator (e.g., StyleGAN [12] and DiffAE [33]) for hybrid domain adaptation, limiting its applicability to a broader range of generators. MTG and DiFa utilize GAN inversion, which restricts the applicability to generators similar to StyleGAN. Conversely, UniHDA is not constrained by the type of generators, allowing for its broader application across various generators.
Figure 11: Ablation of different pre-trained encoders for CSS on hybrid domain adaptation.
$\lambda$ FFHQ (I-I) FFHQ (T-T) FFHQ (T-I) Dog (T-I) Church (T-I)
SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$) CS ($\uparrow$) SCS ($\uparrow$) CS ($\uparrow$) SCS ($\uparrow$) CS ($\uparrow$)
0 0.502 0.639 0.520 0.170 0.562 0.557 0.491 0.430 0.604 0.411
3 0.681 0.638 0.683 0.171 0.694 0.556 0.787 0.428 0.706 0.413
5 0.769 0.642 0.707 0.176 0.742 0.565 0.796 0.430 0.781 0.414
Table 6: Quantitative ablation for our proposed $\mathcal{L}_{\text{CSS}}$.
4.6 Generalization on 3D Generator
To further verify the versatility of UniHDA, we apply it on the popular 3D-aware image generation method, EG3D [4]. Specifically, we replace the discrimination loss with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$ for hybrid domain adaptation. As shown in Fig. 8, the results effectively integrate the attributes and preserve the characters and poses of source domain, demonstrating that UniHDA is agnostic to the type of generative models and can be easily extended to other generators.
4.7 Generalization on Diffusion Model
@@
In this section, we demonstrate that UniHDA is agnostic to the type of generative models and can easily generalize to diffusion models. Specifically, we apply UniHDA on DiffusionCLIP [16] and replace the training objective of DiffusionCLIP with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$. As shown in Fig. 9, the results integrate the characteristics from multiple target domains and maintain robust consistency with the source domain. This verifies desirable generalization ability and versatility of UniHDA. More results are included in Appendix.
4.8 Ablation of CSS Loss
@@
We conduct the ablation study to evaluate the effects of our proposed Cross-domain Spatial Structure (CSS) loss on single domains. As shown in Fig. 10, the results without $\mathcal{L}_{\text{CSS}}$ suffer from overfitting and have very limited cross-domain consistency, e.g., distorted backgrounds in Row 1 and 3. Benefited from CSS, the generated images maintain consistency with the source images in terms of spatial structure, thereby inheriting the diversity from the source domain. More results are included in Appendix. Besides, we conduct the quantitative ablation in Tab. 6, which is consistent with the qualitative results.
4.9 Ablation of Encoder for CSS
We conduct experiments on pre-trained ViT [6], MViTv2 [21], and Dinov2 to explore the impact of different image encoders for CSS. As shown in Fig. 11, we can observe that all of them improve the consistency with source domain compared with the baseline approach. Furthermore, they exhibit a similar qualitative style, which demonstrates that our CSS is agnostic to different pre-trained image encoders. More results are included in Appendix.
@@ 5 Conclusion & Limitation
In this paper, we propose UniHDA, a unified and versatile framework for multi-modal hybrid domain adaptation. To enable multiple modalities, we leverage CLIP encoder to project the references into a unified embedding space. For hybrid domain, we demonstrate the compositional capabilities of direction vectors in CLIP’s embedding space and linearly interpolate direction vectors of multiple target domains. In addition, we propose a new cross-domain spatial structure loss to improve consistency, which is conducted in generator-agnostic space and versatile for various generators. We believe our work is an important step towards generative domain adaptation, since we have demonstrated the source generator can be effectively adapted to a hybrid domain with multi-modal references and maintain robust cross-domain consistency. Our code will be made public.
@@
While UniHDA effectively realizes multi-modal hybrid domain adaptation, it also has the limitation. To encode both image and text into a shared embedding space, we utilize pre-trained CLIP during training time, which might bring potential bias for some domains. Nevertheless, we believe that the exploration of the novel task is significant for future work and solutions could be integrated into UniHDA to eliminate the bias.